A mile is 5280 feet and one kilometer is equal to .62 of mile. How many inches are there in 3.5 kilometers?
Answer by ikleyn(52776)

.

    3.5 kilometers = 3.5*0.62 miles = 3.5*0.62*5280 feet = 3.5*0.62*5280*12 inches = 137,491.2 inches.     ANSWER